A boat goes downstream and covers the distance between two ports in 4 hours while it covers the same distance upstream in 5 hours. If the speed of stream is 2KMPH, what is the speed of boat in still water??

Grade:8

1 Answers

Arun
5 years ago
Let 'x' be the speed of the streamer in still water.
Distance in both cases is same (Suppose it as d).  We know, speed=distance/time.      So when downstream,
x + 2 = d/4
=> d = 4x + 8.  When upstream,
x - 2 = d/5
=> d = 5x - 10.  Since, d is same
=>4x + 8 = 5x - 10
=> 5x - 4x = 10+8
=> x = 18.
Hence, speed of the streamer in still water is 18 kmph.